the perimeter of a triangle is 128 meters. find the legenth of each side if:
side a is: (5x)m
side b is: (6x+1)m
side c is: (6x+8)m

Answer by rfer(16322)

5x+6x+1+6x+8=128
17x+9=128
17x=128-9
x=119/17
x=7
5x=35
8x+1=43
6x+8=50